OpenStudy (anonymous):

Suppose a continuous random variable X is uniformly distributed with 0≤X≤20. What is P(0≤X≤5). Give a decimal answer

OpenStudy (anonymous):

Judging by the given info, it sounds like the density function is \[f(x)=\begin{cases}\dfrac{1}{20}&\text{for }0\le x\le20\\\\0&\text{otherwise}\end{cases}\] The probability is given by the integral, \[P(0\le X\le 5)=\int_0^5f(x)~dx\] Alternatively, if you'd prefer a pre-calculus approach, you can consider the density function graphically:  The probability is the area of the shaded region.